OpenStudy (anonymous):

Equal amount of heat is given to one mole of a gas separately at constant pressure, and at constant volume then, the temperature of the gas will become: * more at constant pressure than at constant volume. * More at constant volume then at constant pressure *Equally in both conditions Which is the correct option?

OpenStudy (anonymous):

Puhlease reply...

OpenStudy (anonymous):

Q = dU + dW at constant V, dT=Q/nC(V) at constant P dT = (Q - PdV)/nC(P)

OpenStudy (anonymous):

it seems the temperature is more at Const. V but i guess more info needs to be supplied

OpenStudy (anonymous):

Yeah please ask someone to help here